OpenStudy (anonymous):

What is the growth factor of the following example? Assume time is measured in the units given. What usage is increasing by 3% per year

OpenStudy (anonymous):

Looks like some part of the question is missing... Im guessing you want the compounding interest function.. and you are after a value like the rate.. P=A(1+r)^(t-1) t = term in this case t=years r= rate of growth as a decimal percentage IE .03 = 3% or 1+r=1.03 A= starting amount